Nine students in Mr. Ruiz's class all took a spelling test. Raw scores can range between 0 (0%) and 10 (100%). Their scores are listed below.

Student   Raw Score
A 4
B 5
C *5*
D 6
E 6
F 6
G 7
H 7
I 8

The class mean is 6, median is 6 , mode is 6 , range is 4 , and standard deviation is 1.1(round to one decimal place).

Look at Student C's score. The score is in the 33 th percentile rank (round to the nearest integer). Its Z score is -0.90 (round to two decimal places, e.g., if you got 0.553, enter .55 or 0.55). Its T score is 41 (round to one decimal place, e.g., if you got 67.38, enter 67.4).

Nine students in Mr. Sotelo's class all took a spelling test. Raw scores can range between 0 (0%) and 10 (100%). Their scores are listed below.

Student   Raw Score
J 2
K 3
L 4
M 4
N *5*
O 6
P 10
Q 10
R 10

The class mean is 6 , median is 5 , mode 10 is , range is 8 , and standard deviation is 3.2 (round to one decimal place).

Look at Student N's score. The score is in the 50 th percentile rank (round to the nearest integer). Its Z score is -0.33(round to two decimal places). Its T score is -0.9 (round to one decimal place).

QUESTION:

What if a 70% correct score was used as a criterion to determine success ? What proportion of students in each class would meet that criterion?

Mr. Ruiz's class: ______ % (Round to the nearest integer)

Mr. Sotelo's class:______ % (Round to the nearest integer)
Answer #1

(1)

Mr. Ruiz's class:

Converting marks out of 10 to percentages, we get:

Student Raw score Percentage score
A 4 40
B 5 50
C 5 50
D 6 60
E 6 60
F 6 60
G 7 70
H 7 70
I 8 80

Given: 70% correct score was used as a criterion to determine success.

So,

Students: A,B,C,D,E and F fail. Students G , H and I pass.

Thus,

Proportion of students would meet that criterion = (3/9)X 100 = 33 %

(2)

Mr. Sotelo's class:

Converting marks out of 10 to percentages, we get:

Student Raw score Percentage score
J 2 20
K 3 30
L 4 40
M 4 40
N 5 50
O 6 60
P 10 100
Q 10 100
R 10 100

Given: 70% correct score was used as a criterion to determine success.

So,

Students: J, K, L, M, N and O fail. Students P , Q and R pass.

Thus,

Proportion of students would meet that criterion = (3/9)X 100 = 33 %

Thus,

Answer is:

Mr. Ruiz's class: 33 %

Mr. Sotelo's class: 33 %